if x + y =12 and xy = 163, find x squared + y squared
Found 2 solutions by edjones, solver91311:
Answer by edjones(8007)

x+y=12
xy=163
.
x=12-y
y(12-y)=163
12y-y^2=163
y^2-12y+163=0
y=6-sqrt(127)i, x=6+sqrt(127)i (or vice versa) [quadratic equation]
x^2+y^2
=[-91+12sqrt(127)i]+[-91-12sqrt(127)i]
=182
.
Ed

Answer by solver91311(24713)



Hence:



Put in standard form and solve:





Therefore:

So:



and




Either way,